Two spherical masses are at a distance d and attract each other with a gravitational force of magnitude P.
What should be the distance d’ between them for the force to become 1/2P
1
Expert's answer
2018-05-28T17:33:08-0400
Newton’s law of universal gravitation states
P=G (m_1 m_2)/d^2
So
1/2 P=G (m_1 m_2)/(2d^2 )=G (m_1 m_2)/d^'2
Therefore
d^'2=2d^2
d^'=√2 d